if xy = 10 and xw = 12, what is vy?
Found 2 solutions by MRperkins, oberobic:
Answer by MRperkins(300)

I think you have something written wrong, we know nothing about v

Answer by oberobic(2304)

x*y = 10
x*w = 12
.
There is no way to tell what "vy" might be.
.
But I assume you're asking, "What is w*y?"
.
The fundamental issue is that there are too many unknowns.
You have 3 unknowns {w, x, y} and only 2 equations.
You need at least as many equations as unknowns to solve the problem.
.
In addition, there are infinitely many ways that x*y = 10.
1*10
2*5
-1*-10
0.01 * 1000
etc.
.
Similarly, there are infinitely many ways that x*w = 12.
.
That means there are infinitely many values for w*y, too.